OpenStudy (anonymous):

What is the area, to the nearest tenth, of a circular rug, whose diameter is 3 feet? Use 3.14 for pi.

jimthompson5910 (jim_thompson5910):

use the formula A = pi*r^2

jimthompson5910 (jim_thompson5910):

in this case, the diameter is 3 feet, so the radius is 1.5 feet (take half of 3 to get 3/2 = 1.5)

OpenStudy (anonymous):

so the answer is 9.4

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

how did you get that?

OpenStudy (anonymous):

3.14 times 1.5 times 2

jimthompson5910 (jim_thompson5910):

I see, r^2 means r squared so you don't multiply by 2, you are squaring

jimthompson5910 (jim_thompson5910):

If you wanted to multiply by 2, you would say r*2 and not r^2

jimthompson5910 (jim_thompson5910):

see the difference?

OpenStudy (anonymous):

yeah im new to this and don't get all the little abbreviations and stuff yet

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

so the answers 4.7

jimthompson5910 (jim_thompson5910):

A = pi*r^2 A = 3.14*1.5^2 A = 3.14*2.25 A = 7.065

jimthompson5910 (jim_thompson5910):

following PEMDAS, you are supposed to square first, then multiply

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the area, to the nearest tenth, of a circular rug, whose diameter is 3 feet? Use 3.14 for pi. A. 4.7 square feet B. 7.1 square feet C. 9.4 square feet D. 18.8 square feet

jimthompson5910 (jim_thompson5910):

you have to round 7.065 to the nearest tenth it looks like

OpenStudy (anonymous):

7.1

jimthompson5910 (jim_thompson5910):

good